OpenStudy (anonymous):

Can you please explain this to me? If f(x) varies directly with x, and f(x) = 12 when x = 8, write the direct linear variation equation. @MajicMuzyk @mathstudent55

OpenStudy (mathstudent55):

We had a problem before that was f(x) = 50x, remember?

OpenStudy (anonymous):

yes I remember

OpenStudy (anonymous):

so f(8)=12?

OpenStudy (mathstudent55):

We found the constant of variation to be 50. In general, y = mx, or f(x) = mx is a direct variation, and m is the constant of variation. You may know m already as the slope.

OpenStudy (anonymous):

right but then what?

OpenStudy (mathstudent55):

If you have a direct variation problem, you know it's going to be f(x) = mx All you need to find now is what value m will have.

OpenStudy (mathstudent55):

For that, you use the given point (8, 12)

OpenStudy (mathstudent55):

Since you are told that when x = 8, f(x) = 12, that means f(8) = 12.

OpenStudy (anonymous):

ok but that is that my answer? because shouldnt it be f(x)=12x or f(x)=8x

OpenStudy (mathstudent55):

Now use the equation I gave you above, replace x with 8 and f(x) with 12, and solve for m. \(f(8) = m \times 8= 12 \) \(8m = 12\) \(m = \dfrac{12}{8} \) \(m = \dfrac{3}{2} \)

OpenStudy (anonymous):

so then it turns into f(x)=3/2?

OpenStudy (mathstudent55):

Almost. You need the x. Now we know that m, the constant of variation is 3/2, we replace m with 3/2: \(f(x) = \dfrac{3}{2}x\) That is the final answer.

OpenStudy (anonymous):

Oh my gosh thank you so much that was stressful! :)

OpenStudy (mathstudent55):

To be sure, try to find f(8). It should be equal to 12: \(f(x) = \dfrac{3}{2} x\) \(f(8) = \dfrac{3}{2} \times 8 \) \(f(8) = \dfrac{3}{2} \times \dfrac{8}{1} \) \(f(8) = \dfrac{24}{2} \) \(f(8) = 12\) This shows the function is correct.

OpenStudy (mathstudent55):

You are welcome.

OpenStudy (anonymous):

ok wow that was a great help!!! thanks again! @mathstudent55